# Sigmoid Head for Quality Estimation under Language Ambiguity

### v1

@@ Abstract
Tu Anh Dinh  and Jan Niehues Karlsruhe Institute of Technology Karlsruhe, Germany {firstname}.{lastname}@kit.edu
@@ 1 Introduction
  • We identify architectural and training setup issues in current LMs that lead to ambiguity-induced underconfidence.
  • We propose the Sigmoid Head  [^1] Implementation available at https://github.com/TuAnh23/sigmoid-confidence. on top of pre-trained LMs for Quality Estimation that addresses these issues. The module is trained using the same data as standard generative training, requires no additional labeled data, and is computationally lightweight during training and inference.
@@ 2.2 Motivation: Issues of Standard LMs
Language ambiguity is learned implicitly by language models. As discussed above, when the training data shows that multiple outputs are correct, the model learns to spread probability mass across these tokens. This behavior was also observed by Dinh and Niehues (2025). They show that, specifically for ambiguous generation tasks, multiple tokens often receive relatively high probability compared to the remaining tokens at each output step. They refer to these tokens as "dominant tokens", which are likely to correspond to valid alternative outputs. They propose a heuristic to identify such tokens. We adopt this idea in our method, described as follows.
@@ 3 Proposed: Sigmoid Head for QE
Unlike softmax, sigmoid removes the constraint that probabilities must sum to one. As a result, multiple tokens can simultaneously receive high probability, which better matches settings where multiple outputs are correct.
At each generation step $i$, we train the Sigmoid Head to distinguish correct from incorrect tokens, given the one-hot encoded target. Let $y_{i}^{\ast}$ denote the single reference token in the one-hot encoded target. The reference token is treated as a positive example with label 1. Negative examples are sampled from the non-reference tokens in the vocabulary $\mathcal{V}\setminus\{y_{i}^{\ast}\}$. We sample 10 negative tokens, as prior work has shown that this is an effective choice for different tasks Mikolov et al. (2013, 2018). Negative tokens are sampled from a configurable distribution. Unless stated otherwise, we use token-frequency-based sampling, where tokens that appear more often in the training data are more likely to be selected as negatives. In this way, a token is expected to be selected as a negative as often as it appears as a positive. This balances the training signal and reduces bias toward frequent tokens. In Section 5.2, we additionally evaluate alternative negative sampling strategies and analyze their impact on performance.
Ambiguity-Informed Negative Sampling
Recall disadvantage D2, where some non-reference tokens from $\mathcal{V}\setminus\{y_{i}^{\ast}\}$ may be valid output alternatives. To reduce the risk of sampling such tokens as negatives, we make use of the observation from Dinh and Niehues (2025): in a trained LM, the tokens with dominant probability mass in the softmax distribution are likely to be the alternative correct options. We therefore apply their heuristic (See Appendix B for details) to identify the set of dominant tokens $\mathcal{D}_{i}$ from the original softmax distribution $P_{\theta}(y_{i}\mid\mathbf{x},\mathbf{y}_{<i})=\mathrm{softmax}(\mathbf{z}^{\text{out}}_{i})$ and exclude them from negative sampling. Our approach can be viewed as a form of implicit knowledge distillation, as we are using the knowledge from the pretrained transformer $\theta_{\text{tr}}$ and softmax head $\theta_{\text{out}}$ to train our new sigmoid head $\theta_{\text{qe}}$. Formally, during training of the sigmoid head, negative samples are drawn from
$\mathcal{N}_{i}=\mathcal{V}\setminus\left(\{y_{i}^{\ast}\}\cup\mathcal{D}_{i}\right),$
@@
The Sigmoid Head is trained with binary cross-entropy loss over the selected samples, consisting of the reference token $y_{i}^{\ast}$ and the negative samples $\mathcal{N}_{i}$. The loss at generation step $i$ is:
@@ 4.1 Models and Training Data
Details of the language models (LMs) used in our paper and their corresponding training data can be found in Table 1. We train Transformer Base from scratch using the base configuration of Fairseq Ott et al. (2019) on 5M samples of ParaCrawl English-German Machine Translation data, filtered by Bicleaner AI Zaragoza-Bernabeu et al. (2022); de Gibert et al. (2024). DeltaLM is finetuned on the same ParaCrawl data. Tower Instruct v2 and OLMo SFT v2 models are taken off the shelves. We experiment with adding our Sigmoid Head for Quality Estimation onto these models. More details on our motivation for model choices are in Appendix C.
@@ 4.2 Test Data and Evaluation
We evaluate the Quality Estimation performance of our Sigmoid Head on three different generation tasks: Machine Translation, Paraphrase, and Question Answering. We place more focus on the Machine Translation (MT) task, as MT evaluation is more well-studied, with different established data, models, and benchmarks thanks to the WMT Shared Tasks Kocmi et al. (2024); Freitag et al. (2024); Zerva et al. (2024).
@@
More details on the test sets and choice of language pairs can be found in Appendix D.
@@ 5.1 Overall Performance
Our approach also works on providing QE scores on the word level, allowing DeltaLM to outperform supervised QE. Details are in Appendix E.
@@
Table 5: Comparison to other common QE approaches.
@@ 5.2 Effect of Negative Sampling Strategies
Avoiding Dominant Tokens
@@ 7 Conclusion
The main motivation of this work is to improve the training setup and model architecture to better account for the ambiguity of natural language. In standard large language model training, such ambiguity already exists during next-token prediction pretraining, where multiple tokens can be valid at each generation step. However, due to limited computational resources, we only apply our proposed training setup to the final instruction-following training stage. Applying our approach during the pretraining phase could potentially bring stronger benefits, as the model could learn ambiguity more effectively from the large-scale pretraining data. Another limitation is that we evaluate our approach on only two LLMs, Tower Instruct 7B and OLMo 1B, as these are among the few model families for which the training data is publicly available.
@@ Acknowledgements
This work was supported by the Helmholtz Programme-oriented Funding, with project number 46.24.01, project name AI for Language Technologies. We acknowledge the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Wurttemberg and by the Federal Ministry of Education and Research.
@@ Appendix C Motivation for Model Choices
Appendix D Test Sets for Evaluation
@@
Appendix E Word-level QE
@@
Appendix F Hardware
@@
Appendix G License For Artifacts

### v2

@@ Abstract
Tu Anh Dinh  and Jan Niehues Karlsruhe Institute of Technology Karlsruhe, Germany tu.dinh@kit.edu, jan.niehues@kit.edu
@@ 1 Introduction
  • We identify LMs’ architecture and training issues causing ambiguity-induced underconfidence.
  • We propose a Sigmoid Head  [^1] Implementation available at https://github.com/TuAnh23/sigmoid-head-qe. on top of pre-trained LMs for Quality Estimation to address these issues. The module is trained on the same data as standard generative training, requires no additional labeled data, and is computationally lightweight during training and inference.
@@ 2.2 Motivation: Issues of Standard LMs
Language ambiguity is learned implicitly by language models. As discussed above, when the training data shows that multiple outputs are correct, the model learns to spread probability mass across these tokens. This behavior was also observed by Dinh and Niehues (2025). They show that, specifically for ambiguous generation tasks, multiple tokens in the softmax distribution often receive relatively high probability compared to the remaining tokens at each output step. They refer to these tokens as "dominant tokens", which are likely to correspond to valid alternative outputs. They propose a heuristic to identify such tokens. We adopt this idea in our method, described in Section 3.
2.3 Noise-Contrastive Estimation
An alternative to the final softmax activation in standard LM training is to use the sigmoid function, as proposed by Mnih and Teh (2012), known as Noise-Contrastive Estimation (NCE). NCE reformulates next-token training as a binary discrimination task: the reference token is treated as a positive example, while several tokens sampled from a predefined noise distribution are treated as negatives. The model is trained to assign high probability to the reference token and low probability to the sampled noise tokens using a sigmoid-based objective. The goal is to avoid the need for full softmax calculation during training and reduces computation for large vocabularies.
As the softmax function may not be well suited to model the ambiguity of language (Section 2.2), we make use of NCE in our method, described in Section 3.
@@ 3 Proposed: Sigmoid Head for QE
Unlike softmax, sigmoid removes the constraint that probabilities must sum to one. Multiple tokens can simultaneously receive high probability, better matching settings where multiple outputs are correct.
At each output step $i$, we train the Sigmoid Head to distinguish correct from incorrect tokens, given the one-hot encoded target. Let $y_{i}^{\ast}$ denote the single reference token in the one-hot encoded target. The reference token is treated as a positive example with label 1. Negative examples are sampled from the non-reference tokens in the vocabulary $\mathcal{V}\setminus\{y_{i}^{\ast}\}$. We sample 10 negative tokens, as prior work has shown that this is an effective choice for different tasks Mikolov et al. (2013, 2018). Negative tokens are sampled from a configurable distribution. Unless stated otherwise, we use token-frequency-based sampling, where tokens that appear more often in the training data are more likely to be selected as negatives. In this way, a token is expected to be selected as a negative as often as it appears as a positive. This balances the training signal and reduces bias toward frequent tokens. In Section 5.2, we evaluate alternative negative sampling strategies and analyze their impact on performance.
Ambiguity-Informed Negative Sampling: Avoid Dominant Tokens
Recall disadvantage D2, where some non-reference tokens from $\mathcal{V}\setminus\{y_{i}^{\ast}\}$ may be valid output alternatives. To reduce the risk of sampling such tokens as negatives, we make use of the observation from Dinh and Niehues (2025): in a trained LM, the tokens with dominant probability mass in the softmax distribution are likely to be the alternative correct options. We therefore apply their heuristic (See Appendix B for details) to identify the set of dominant tokens $\mathcal{D}_{i}$ from the original softmax distribution $P_{\theta}(y_{i}\mid\mathbf{x},\mathbf{y}_{<i})=\mathrm{softmax}(\mathbf{z}^{\text{out}}_{i})$ and exclude them from negative sampling. Our approach can be viewed as a form of implicit knowledge distillation, as we are using the knowledge from the pretrained transformer $\theta_{\text{tr}}$ and softmax head $\theta_{\text{out}}$ to train our new sigmoid head $\theta_{\text{qe}}$. Formally, during training of the sigmoid head, negative samples $\mathcal{N}_{i}$ are drawn from
$\mathcal{V}\setminus\left(\{y_{i}^{\ast}\}\cup\mathcal{D}_{i}\right),$
@@
The Sigmoid Head is trained with binary cross-entropy loss over the selected samples, consisting of the reference token $y_{i}^{\ast}$ and the negative samples $\mathcal{N}_{i}$. The loss at output step $i$ is:
@@
During inference, the output text is generated by the original LM components $\theta=\{\theta_{\text{tr}},\theta_{\text{out}}\}$ without modification. The Sigmoid Head $\theta_{\text{qe}}$ operates in parallel by taking the last hidden states from $\theta_{\text{tr}}$ at each generation step and produces a quality score for each output token.
@@ 4.1 Models and Training Data
Details of the language models (LMs) used in our paper and their corresponding training data can be found in Table 1. We train Transformer Base from scratch using the base configuration of Fairseq Ott et al. (2019) on 5M samples of ParaCrawl English-German Machine Translation data, filtered by Bicleaner AI Zaragoza-Bernabeu et al. (2022); de Gibert et al. (2024). DeltaLM is finetuned on the same ParaCrawl data. Tower Instruct v2 and OLMo SFT v2 models are taken off the shelves. We experiment with adding our Sigmoid Head for Quality Estimation onto these models. More details on our motivation for model choices are in Appendix C. We discuss the increase in model parameter counts after adding the Sigmoid Head in Appendix D.
@@ 4.2 Test Data and Evaluation
We evaluate the Quality Estimation performance of our Sigmoid Head on three different generation tasks: Machine Translation, Paraphrase, and Question Answering. The QE score is obtained for each model output by multiplying the token-level scores produced by the Sigmoid Head. We place more focus on the Machine Translation (MT) task, as MT evaluation is more well-studied, with different established data, models, and benchmarks thanks to the WMT Shared Tasks Kocmi et al. (2024); Freitag et al. (2024); Zerva et al. (2024).
@@
More details on the test sets and choice of language pairs can be found in Appendix E.
@@ 5.1 Overall Performance
Our approach also works on providing QE scores on the word level, allowing DeltaLM to outperform supervised QE. Details are in Appendix F.
@@
We provide a qualitative analysis in Appendix G, presenting example cases where the original Softmax Head fails but the Sigmoid Head succeeds.
@@
Table 5: Comparison to common QE approaches.
@@ 5.2 Effect of Negative Sampling Strategies
Ambiguity-Informed Negative Sampling: Avoid Dominant Tokens
@@ 7 Conclusion
The main motivation of this work is to improve the training setup and model architecture to better account for the ambiguity of natural language. In standard large language model training, such ambiguity already exists during next-token prediction pretraining, where multiple tokens can be valid at each generation step. However, due to limited computational resources, we only apply our proposed training setup to the final instruction-following training stage. Applying our approach during the pretraining phase could potentially bring stronger benefits, as the model could learn ambiguity more effectively from the large-scale pretraining data. Another limitation is that we evaluate our approach on only two LLMs, Tower Instruct 7B and OLMo 1B, as these are among the few model families for which the training data is publicly available. Lastly, adding the Sigmoid Head increases the parameter count of language models, though the resulting overhead is relatively minor for most state-of-the-art large models.
@@ Acknowledgements
This work was supported by the Helmholtz Programme-oriented Funding, with project number 46.24.01, project name AI for Language Technologies. We acknowledge the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Wurttemberg and by the Federal Ministry of Education and Research. This work also received support from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People).
@@ Appendix C Motivation for Model Choices
Appendix D Increase in Parameter Count with Sigmoid Head
We report parameter counts before and after adding the Sigmoid Head in Table 7. Introducing an additional unembedding head increases memory usage. This overhead is more pronounced for models with relatively small total parameter counts but large vocabularies (e.g., DeltaLM and OLMo 1B). However, for many state-of-the-art large models, the unembedding layer constitutes only a small fraction of the total parameters, for example, 5.87% for OLMo-3-7B, 2.77% for gpt-oss-20B, 0.496% for gpt-oss-120B, and 0.265% for Qwen3-235B-A22B. In such cases, the additional memory cost of the Sigmoid Head is relatively minor.
Model Base Model Base + Increase
Size Sigmoid Head (%)
Transformer Base 0.062B 0.068B 9.7%
DeltaLM 0.830B 1.086B 30.8%
Tower Instruct v2 7.000B 7.131B 1.9%
OLMo SFT v2 1.000B 1.206B 20.6%
Table 7: Model parameter count before and after adding the Sigmoid Head.
Appendix E Test Sets for Evaluation
@@
Appendix F Word-level QE
@@
Appendix G Qualitative Analysis: When Does Softmax Fail?
Input sentence Gold translation MT output Human Sigmoid Softmax Softmax scores of
score score score valid alternatives
Kurze Notiz über die Presse – der Freitag Short note about the press – Friday Brief note on the press – Friday 0.98 0.99 0.02 Short 0.40, A 0.34, Brief 0.19, Quick 0.04
Transgender: Identifiziert sich nicht mit dem bei der Geburt zugewiesenen Geschlecht. Transgender: Does not identify with the gender assigned at birth. Transgender: Does not identify with the sex assigned at birth. 0.99 0.98 0.02 gender 0.74, sex 0.20
Haushaltspolitik: Steuereinnahmen niedriger als erwartet Budget policy: Tax revenues lower than expected Budgetary policy: Tax revenue lower than expected 0.92 0.98 0.00 Budget 0.52, Budgetary 0.01
Funktion: Weiterschalten auf nächsten Programmschritt (Anzeige) Function: Advance to next programming step (display) Function: Switch to the next program step (display) 0.97 0.99 0.00 Proceed 0.12, Continue 0.10, Move 0.10, Go 0.09, Advance 0.08, Switch 0.05
Schnelle Lieferung und gute Verarbeitung Fast delivery and good manufacturing Fast delivery and good workmanship 0.96 0.98 0.19 processing 0.57, workmanship 0.30, handling 0.03, crafting 0.02, manufacturing 0.02
Ekelhafter, dauerhafter Geruch Disgusting, permanent smell Disgusting, persistent odor 1.00 0.99 0.07 persistent 0.35, permanent 0.25, lasting 0.12; smell 0.51, odor 0.03, odour 0.01
Table 10: Example of good MT outputs along with their QE scores from human, Sigmoid Head, and Softmax Head.
We provide example sentences in Table 10 where the Softmax Head fails while the Sigmoid Head succeeds. Through manual inspection, we identify many cases where the translations are adequate, but the Softmax Head assigns very low scores because one or more tokens have multiple valid alternatives. We highlight such tokens in bold. In the last column, we report the softmax probabilities of the top valid alternatives (word-level scores), showing that these tokens compete for probability mass under softmax, leading to ambiguity-induced underconfidence. The data used is WMT23 German–English, and the scoring model is Tower Instruct v2.
Appendix H Hardware
@@
Appendix I License For Artifacts